OpenStudy (anonymous):

X-y+z=5 2y+3z=14 -3y+2z=5

OpenStudy (raden):

question ?

OpenStudy (anonymous):

How to find y and z ?

OpenStudy (raden):

use elimination and subtitution method. 2st equation times 2, we get : 4y+6z=28 2rd equation times 3, we get : -9y+6z=15 subtract both equations above : 4y+6z=28 -9y+6z=15 ----------- - 13y = 13 divideb both sides by 13 we get 13y/13 = 13/13 y = 1 now subtitute y = 1 to one of equations above (let to 2nd eqution) 2y+3z=14 2(1) + 3z = 14 2 + 3z = 14 subtracted by 2 on both sides, we get 2 + 3z - 2 = 14 - 2 3z = 12 divided by 3 both sides, we get 3z/3 = 12/3 z = 4

OpenStudy (raden):

# divided